# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v1

@@ 4.4.1 In-depth Analysis of Attention Mechanism
To better understand the mechanism of our model, which focuses on the most discriminative information in the brain for the diagnosis of brain diseases, we first performed a visualization analysis of the attention scores $\gamma_{1}$, $\gamma_{2}$ of the attention-guided feature fusion mechanism for all patients of each type of disease, using ADNI and ABIDE as examples. Fig. 6 shows that our model is capable of adaptively learning weights according to the characteristics of different diseases, thereby integrating higher- and lower-order topological information for the diagnosis of brain diseases. In addition, we observed that quadruplet-level interaction signatures are significantly more important than two-dimensional void descriptors and edge features, with their average attention scores exceeding 0.5 on both datasets.
Next, we also conducted a visualization analysis of the attention scores $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$ derived from the signed high-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Fig. 7 shows that across both datasets, the importance of negative synergistic quadruplets is generally greater than that of positive synergistic information in most cases, while the importance of positive synergistic voids exceeds that of negative synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
To further investigate the aforementioned phenomena in the attention mechanism, we conducted inter-group comparisons of the signed high-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors, captured by the model in both patient and healthy control groups, using ADNI dataset as example. As shown in Fig. 8, we observed that the group differences in the quadruplet-level interaction signatures were highly significant, with all statistical significance levels at $p<0.001$ compared to void descriptors. This finding aligns with the results of our model visualization. This may be because, during disease progression, the direct disruption occurs in the higher-order interactions within the brain, which further affects the high-dimensional hole structures in the brain. Moreover, as the disease progresses, the number of positive quadruplets shows a gradual decreasing trend, while the number of negative quadruplets first decreases and then increases. This indicates that during the transition from CN (cognitively normal) to MCI (mild cognitive impairment), the synergistic interactions in the brain—whether positive or negative—decrease. In contrast, as the condition progresses from MCI to AD (Alzheimer’s disease), brain regions tend to act negative-synergistically and simultaneously reduce their excitability in order to maintain the functionality of various neural modules. Regarding the number of voids, although no clear trend was observed in the progression of the disease, we found that positive voids exhibited more significant intergroup differences compared to negative voids. This further supports the earlier finding that they have higher attention scores. Accordingly, CN tends to utilize positive information to construct the whole-brain topological structure, whereas AD appears to rely more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@
Figure 7: The visualization of attention scores for quadruplet in (a) the ADNI dataset and (b) the ABIDE dataset, and the visualization of attention scores for void in (c) the ADNI dataset and (d) the ABIDE dataset
@@
Figure 8: Group difference comparison of the signed high-order topological feature quadruplet-level interaction signatures and two-dimensional void descriptors on ADNI. Statistical significance is denoted as follows: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$, and ns $p\geq 0.05$.”
@@ 4.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Fig. 9, 10. The attention scores are averaged across all subjects in the ADNI and ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both low-order and high-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrixs on ADNI and ABIDE dataset in Fig. 11. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and high-order patterns, which further demonstrates that varying degrees of modularity in both low-order and high-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.
(a) three cluster assignment matrixs on ADNI.
(b) three cluster assignment matrixs on ABIDE
Figure 11: Visualization of three cluster assignment matrixs on ADNI and ABIDE datasets
@@ 4.4.3 The important brain regions and interactions of brain regions
In this section, we further investigate whether the learned ROIs attention scores are interpretable and consistent with previous findings. Here, we take the ADNI, PPMI and ABIDE datasets as examples for discussion of three dieases. These three datasets contain larger sample sizes, enabling the reflection of more generalizable phenomena for brain diseases. First, we analyze the important brain regions and interactions of brain regions associated with brain diseases using HOI-Brain from a more comprehensive perspective rather than a low-level perspective as in previous studies (Kan et al., 2022b; Yu et al., 2024). For each subject, HOI-brain gets three attention matrixes at the first layer of multi-head self attention module. Higher attention scores indicate greater discriminative power of the regions for AD, PD, and ASD. For each disease, we obtain a comprehensive attention matrix that simultaneously captures both low-level and high-level patterns by performing a weighted summation of the three attention matrices for each individual, with the weights derived from the attention scores of each channel. At the regional and regional interaction level, we average the integrated attention matrices across all individuals to derive a matrix representing the importance of interactions between regions. At the regional level, we further compute the row-wise mean of this matrix to obtain the final regional importance scores. We visualize the top ten ROIs and interactions of ROIs with the highest attention scores using BrainNet Viewer (Xia et al., 2013), respectively.
@@
Parkinson’s syndrome (PD). As illustrated in Fig. 13, the ten regions of interest (ROIs) with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in Parkinson’s disease (PD). Specifically, the Precentral Gyrus (PCG.R), a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex (ORBsupmed.L) is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with healthy controls (HCs), patients with Parkinson’s disease (PD) showed a globally reduced structural - functional connectivity decoupling in The Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala (AMYG.L), central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus (THA.R), a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus (STG.R) and Parkinson’s disease, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Crucially, nlike ADNI and ASD, the ”core hubs” in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—did not rank among the top positions in the ”key brain importance” list. Conversely, the seven regions listed as ”key brain importance,” such as ACG.R, TPOmid.R, and CAU.L, were almost entirely absent from the top 20 strongest interactions. Together, these findings highlight two distinct perspectives on ”importance” in brain: ”global centrality” versus ”local strong coupling.”
@@
4.4.4 The high-order organization that differ between patients and healthy individuals
@@
Next, we further investigate whether the higher-order interactions among key brain regions identified by HOI-Brain show significant differences during disease progression, which could offer additional therapeutic insights for brain disease research. For ease of calculation, we selected four out of the top ten key brain regions identified for each disease. Using our proposed time-derivative-multiplication-based co-fluctuation calculation method, we computed the quadruplet interactions among the relevant brain regions and subsequently compared the group differences. As shown in Fig. 15, 16, 17, compared to other brain regions, although these brain regions all demonstrate significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD)—indicating that they may have sustained varying degrees of damage—their higher-order interactions exhibit notable differences across various stages of disease progression. This suggests a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves, which is consistent with the literature (Wang et al., 2019). In addition, we observed that during the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD), the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. On the other hand, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s syndrome, which may be due to the fact that Parkinson’s patients, because of their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s syndrome may result from excessive influence of higher-order brain regions on early sensory processing, which is consistent with the enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@
Alzheimer’s disease (AD). As shown in Fig. 15, in particular, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, exhibit significant inter-group differences during the progression from cognitively normal (CN) to mild cognitive impairment (MCI), with p-values <0.05<0.05< 0.05. However, these differences become less pronounced during the progression from MCI to AD, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests that these higher-order interactions may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease. In contrast, the quadruplet interactions among CAU.R, TPOmid.L, HIP.L, and AMYG.L, which involves the ’social-reward-emotion-memory’ integrated circuit, shows a significant inter-group differences throughout the entire AD progression, with p-values <0.001<0.001< 0.001. This indicates that these higher-order interactions may be persistently disrupted across the entire disease course, playing a critical role in AD pathogenesis.
@@
Parkinson’s syndrome (PD). As shown in Fig. 16, the quadruplet interactions among PCG.R, ORBsupmed.L, IPL.R, and SPG.R; the interactions among PCG.R, ORBsupmed.L, IPL.R, and IFGoperc.R; and those among PCG.R, ORBsupmed.L, IPL.R, and STG.R, which collectively engage the dorsal attention network (DAN) and fronto-parietal executive control circuits, exhibit significant inter-group differences ($p<0.05$) between healthy controls and early-stage PD patients. These findings suggest that these functional connectivity patterns may serve as potential early biomarkers for cognitive-motor integration deficits characteristic of the prodromal stages of PD. In addition, the quadruplet interactions involving PCG.R, ORBsupmed.L, THA.R, and IFGoperc.R, which map onto the mesocortical limbic-striatal-thalamic loop, show inter-group differences ($p<0.001$). This persistent disruption aligns with the progressive degeneration of dopaminergic pathways and thalamocortical relay integrity, implicating these interactions in both motor dysfunction (e.g., bradykinesia) and non-motor symptoms (e.g., apathy, executive decline) throughout PD pathogenesis.critical role in PD-related motor and cognitive dysfunction.

### v2

@@ 4.4.1 In-depth Analysis of Attention Mechanism
To better understand the mechanism of our model, which focuses on the most discriminative information in the brain for the diagnosis of brain diseases, we first performed a visualization analysis of the attention scores $\gamma_{1}$, $\gamma_{2}$ of the attention-guided feature fusion mechanism for all patients of each type of disease, using ADNI and ABIDE as examples. Fig. 6 shows that our model is capable of adaptively learning weights according to the characteristics of different diseases, thereby integrating higher-order and lower-order topological information for the diagnosis of brain diseases. In addition, we observed that quadruplet-level interaction signatures are significantly more important than two-dimensional void descriptors and edge features, with their average attention scores exceeding 0.5 on both datasets.
Next, we also conducted a visualization analysis of the attention scores $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$ derived from the signed higher-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Fig. 7 shows that across both datasets, the importance of negative synergistic quadruplets is generally greater than that of positive synergistic information in most cases, while the importance of positive synergistic voids exceeds that of negative synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
In order to conduct a more in-depth investigation into the aforementioned phenomena in the attention mechanism, inter-group comparisons were conducted of the signed higher-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors. These were captured by the model in both the patient and healthy control groups. The ADNI dataset was used as an exemplar. As demonstrated in Fig. 8, which presents a comparison of group differences in higher-order topological features, it was observed that the group differences in the quadruplet-level interaction signatures were found to be highly significant, with all statistical significance levels at $p<0.001$ compared to void descriptors. This finding is consistent with the results of our model visualisation. This phenomenon may be attributed to the disruption of higher-order interactions within the brain during the progression of the disease, which in turn exerts an influence on the high-dimensional hole structures present within the brain. Furthermore, as the disease progresses, the number of positive quadruplets exhibits a gradual decrease, while the number of negative quadruplets initially decreases and subsequently increases. This finding suggests that during the transition from CN (cognitively normal) to MCI (mild cognitive impairment), there is a decline in synergistic interactions in the brain, whether positive or negative. Conversely, as the condition progresses from MCI to AD (Alzheimer’s disease), brain regions have been observed to exhibit negatively synergistic interactions, concurrently reducing their excitability to preserve the functionality of diverse neural modules. With regard to the number of voids, although no clear trend was observed in the progression of the disease, it was found that positive voids exhibited more significant intergroup differences compared to negative voids. This finding lends further support to the hypothesis that they exhibit higher levels of attention scores. Consequently, CN appears to utilise positive information to construct the whole-brain topological structure, whereas AD relies more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@
Figure 7: The visualization of attention scores for quadruplet in (a) the ADNI dataset and (b) the ABIDE dataset, and the visualization of attention scores for void in (c) the ADNI dataset and (d) the ABIDE dataset.
@@
Figure 8: Group difference comparison of the signed higher-order topological feature quadruplet-level interaction signatures and two-dimensional void descriptors on ADNI. Statistical significance is denoted as follows: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$, and ns $p\geq 0.05$.”
@@ 4.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Fig. 9, 10. The attention scores are averaged across all subjects in the ADNI and ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both low-order and higher-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrices on the ADNI and ABIDE datasets in Fig. 11. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and higher-order patterns, which further demonstrates that varying degrees of modularity in both low-order and higher-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.
(a) three cluster assignment matrices on ADNI.
(b) three cluster assignment matrices on ABIDE.
Figure 11: Visualization of three cluster assignment matrices on ADNI and ABIDE datasets.
@@ 4.4.3 The important brain regions and interactions of brain regions
In this section, we further investigate whether the learned ROIs attention scores are interpretable and consistent with previous findings. Here, we take the ADNI, PPMI and ABIDE datasets as examples for discussion of three diseases. These three datasets contain larger sample sizes, enabling the reflection of more generalizable phenomena for brain diseases. First, we analyze the important brain regions and interactions of brain regions associated with brain diseases using HOI-Brain from a more comprehensive perspective rather than a low-level perspective as in previous studies (Kan et al., 2022b; Yu et al., 2024). For each subject, HOI-brain gets three attention matrixes at the first layer of multi-head self attention module. Higher attention scores indicate greater discriminative power of the regions for AD, PD, and ASD. For each disease, we obtain a comprehensive attention matrix that simultaneously captures both low-level and high-level patterns by performing a weighted summation of the three attention matrices for each individual, with the weights derived from the attention scores of each channel. At the regional and regional interaction level, we average the integrated attention matrices across all individuals to derive a matrix representing the importance of interactions between regions. At the regional level, we further compute the row-wise mean of this matrix to obtain the final regional importance scores. We visualize the top ten ROIs and interactions of ROIs with the highest attention scores using BrainNet Viewer (Xia et al., 2013), respectively.
@@
Parkinson’s syndrome (PD). As illustrated in Fig. 13, the ten regions of interest (ROIs) with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in Parkinson’s disease (PD). Specifically, the Precentral Gyrus (PCG.R), a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex (ORBsupmed.L) is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with healthy controls (HCs), patients with Parkinson’s disease (PD) showed a globally reduced structural-functional connectivity decoupling in the Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala (AMYG.L), central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus (THA.R), a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus (STG.R) and Parkinson’s disease, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Crucially, unlike ADNI and ASD, the ”core hubs” in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—did not rank among the top positions in the ”key brain importance” list. Conversely, the seven regions listed as ”key brain importance,” such as ACG.R, TPOmid.R, and CAU.L, were almost entirely absent from the top 20 strongest interactions. Together, these findings highlight two distinct perspectives on ”importance” in brain: ”global centrality” versus ”local strong coupling.”
@@
4.4.4 The higher-order organization that differ between patients and healthy individuals
@@
Next, we further investigate whether the higher-order interactions among key brain regions identified by HOI-Brain show significant differences during disease progression, which could offer additional therapeutic insights for brain disease research. For ease of calculation, we selected four out of the top ten key brain regions identified for each disease. Using our proposed Multiplication of Temporal Derivatives (MTD)-based co-fluctuation calculation method, we computed the quadruplet interactions among the relevant brain regions and subsequently compared the group differences. As shown in Fig. 15, 16, 17, compared to other brain regions, although these brain regions all demonstrate significant importance in the diagnosis of Alzheimer’s disease (AD) or Autism Spectrum Disorder (ASD)—indicating that they may have sustained varying degrees of damage—their higher-order interactions exhibit notable differences across various stages of disease progression. This suggests a dissociation between higher-order interactions in the brain and the functions of the individual brain regions themselves, which is consistent with the literature (Wang et al., 2019). In addition, we observed that during the transition from healthy individuals to those with Alzheimer’s disease (AD) or autism spectrum disorder (ASD), the positive higher-order interactions between relevant brain regions gradually weaken, indicating a progressive impairment of higher-order synergistic functions among brain regions. On the other hand, the negative higher-order interactions between these brain regions gradually strengthen, reflecting an increasing degree of internal functional disruption in the brain. This phenomenon is consistent with findings reported in previous literature (Santoro et al., 2024). However, this phenomenon is exactly the opposite in Parkinson’s disease, which may be due to the fact that Parkinson’s patients, because of their motor impairments, are prone to hallucinations. Hallucinations in Parkinson’s disease may result from excessive influence of higher-order brain regions on early sensory processing, which is consistent with the enhanced functional integration between sensory and higher-order networks (Tan et al., 2023).
@@
Alzheimer’s disease (AD). As shown in Fig. 15, in particular, the quadruplet interactions among CAU.R, HIP.L, PHG.R, and AMYG.L; the quadruplet interactions among OLF.L, HIP.L, PHG.R, and AMYG.L; and the quadruplet interactions among CAU.R, OLF.L, PHG.R, and AMYG.L, which involve the cortico-striatal-olfactory circuit, exhibit significant inter-group differences during the progression from cognitively normal (CN) to mild cognitive impairment (MCI), with p-values <0.05<0.05< 0.05. However, these differences become less pronounced during the progression from MCI to AD, with p-values 0.05$\geq$ 0.05≥ 0.05. This suggests that these higher-order interactions may serve as potential biomarkers for the early diagnosis of Alzheimer’s disease. In contrast, the quadruplet interactions among CAU.R, TPOmid.L, HIP.L, and AMYG.L, which involves the ’social-reward-emotion-memory’ integrated circuit, shows significant inter-group differences throughout the entire AD progression, with p-values <0.001<0.001< 0.001. This indicates that these higher-order interactions may be persistently disrupted across the entire disease course, playing a critical role in AD pathogenesis.
@@
Parkinson’s disease (PD). As shown in Fig. 16, the quadruplet interactions among PCG.R, ORBsupmed.L, IPL.R, and SPG.R; the interactions among PCG.R, ORBsupmed.L, IPL.R, and IFGoperc.R; and those among PCG.R, ORBsupmed.L, IPL.R, and STG.R, which collectively engage the dorsal attention network (DAN) and fronto-parietal executive control circuits, exhibit significant inter-group differences ($p<0.05$) between healthy controls and early-stage PD patients. These findings suggest that these functional connectivity patterns may serve as potential early biomarkers for cognitive-motor integration deficits characteristic of the prodromal stages of PD. In addition, the quadruplet interactions involving PCG.R, ORBsupmed.L, THA.R, and IFGoperc.R, which map onto the mesocortical limbic-striatal-thalamic loop, show inter-group differences ($p<0.001$). This persistent disruption aligns with the progressive degeneration of dopaminergic pathways and thalamocortical relay integrity, implicating these interactions in both motor dysfunction (e.g., bradykinesia) and nonmotor symptoms (e.g., apathy, executive dysfunction) throughout the pathogenesis of PD-related motor and cognitive impairment.